OpenStudy (anonymous):

Combination factoring Common factor and difference of squares of following equation 2x^2-32

OpenStudy (anonymous):

first rewrite as \[2(x^2-16)\] and then factor as the difference of two squares

OpenStudy (anonymous):

how would you write that as the difference of two squares

OpenStudy (anonymous):

\[x^2\] is the square of \[x\] and \[16\] is the square of \[4\] so it is like having \[a^2-b^2\] with \[a=x,b=4\]

OpenStudy (anonymous):

so if for example the equation was 18x^2-32y^2.... how would you figure this out

OpenStudy (anonymous):

you would first change it to 2(9x^2 - 16y^2) ... then what

OpenStudy (anonymous):

you would recognize \[9x^2=(3x)^2, 16y^2=(4y)^2\] and then write \[(3x)^2-(4y)^2=(3x+4y)(3x-4y)\]